OpenStudy (anonymous):

(a+bi)^2=i what a and b satisfies the equation?

OpenStudy (anonymous):

square it first

jimthompson5910 (jim_thompson5910):

(a+bi)^2=i (a+bi)(a+bi)=i a^2+2abi +b^2i^2 = i a^2+2abi +b^2(-1) = i a^2+2abi - b^2 = i (a^2-b^2) + 2ab*i = 0 + 1*i So this means a^2 - b^2 = 0 2ab = 1

OpenStudy (anonymous):

what u can do is expand the left then simplify... then match up the real parts and the imaginary parts...

jimthompson5910 (jim_thompson5910):

You now have a system of two equations with 2 unknowns. So you can solve for both 'a' and 'b' using this system.

OpenStudy (anonymous):

could you please solve I do not understand fully

jimthompson5910 (jim_thompson5910):

Solve the second equation for any variable (I'm choosing b) 2ab = 1 b = 1/(2a) Now plug this into the other equation a^2 - b^2 = 0 a^2 - (1/(2a))^2 = 0 a^2 - 1/(4a^2) = 0 4a^4 - 1 = 0 I'll let you finish

OpenStudy (phi):

another way is express i in polar form \[ i= e^{i \frac{\pi}{2}} \] if you want sqrt(i): \[ i^{\frac{1}{2}}= (e^{i \frac{\pi}{2}} )^\frac{1}{2}=e^{i \frac{\pi}{4}} \] or cos(pi/4) + i sin(pi/4) sqrt(2)/2 + i sqrt(2)/2

OpenStudy (anonymous):

phi could you explain your way?

OpenStudy (phi):

do you know e^ix = cos(x) + i*sin(x) ? if x= pi/2, cos(pi/2)=0 and sin(pi/2)= 1 so e^(i*pi/2) = i to take the square root, raise e^(i*pi/2) to the 1/2 power (multiply exponents) change back to "rectangular form" using e^ix = cos(x) + i*sin(x)

OpenStudy (anonymous):

so a+bi is just another way of writing cos(x) + i*sin (x)?

OpenStudy (phi):

almost. a+bi --> sqrt(a^2+b^2) * e^(atan(b/a)) See http://en.wikipedia.org/wiki/Complex_number#Polar_form it is helpful to know the magnitude of i is 1.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (phi):

If you don't know this, then stay with rectangular form and solve for a and b as jt showed.